The rms current through a light bulb is 0.718 A when the bulb is connected to a standard 120 V rms source.

(a) Find the bulb's resistance.


(b) Find the peak instantaneous current.


(c) Find the peak instantaneous voltage.


(d) Find the average electrical power used by the bulb.
Answer #1

Vrms= 120 V , Irms= 0.718 A

a) Vrms= Irms*R

R= ( Vrms/ Irms)= 120/0.718= 167.13 ohm.

b) Peak current Io= √2 Irms =√2 * 0.718 = 1.02 A

c) Peak voltage Vo=√2 Vrms= √2 * 120= 169.68 V

d) Average Power in a Cycle, Pav= Vrms * Irms= 120*0.718= 86.16 W